# The Computational Limits of State-Space Models and Mamba via the Lens of Circuit Complexity

### v1

@@ Appendix D More Related Work
Attention optimization.
Besides Mamba, various techniques have been developed to optimize the approximation of attention computation in the transformer architecture, aiming to address the quadratic complexity. Both Mamba and Transformer are LLMs. Attention optimization includes methods for optimizing attention-related regression problems [5, 6, 65, 27, 29, 28, 46, 43, 18, 45, 64, 26, 42, 17, 16, 71], multi-layer attention optimization [63, 52, 49], cross-attention mechanisms [50], applications of Hopfield Models [39, 69, 35, 72, 68, 33, 34, 38], and approaches to enhance the tensor-based attention approximation [51, 7].

### v2

@@ Appendix D More Related Work
Theoretical Machine Learning.
Our work also takes inspiration from the following Machine Learning Theory work. Some works analyze the expressiveness of a neural network using the theory of complexity [72, 60, 67, 25, 66, 55, 51]. Some works optimize the algorithms that can accelerate the training of a neural network [75, 62, 30, 32, 127, 126, 31, 23, 111, 109, 81, 89, 53, 54, 52, 15, 33, 113, 42, 43, 44, 40, 116, 87, 90, 51, 48]. Some works analyze neural networks via regressions [24, 38, 91, 41, 105, 27, 112, 115, 108, 114, 71]. Some works use reinforcement learning to optimize the neural networks [123, 124, 92, 93, 79, 77, 88]. Some works optimize the attention mechanisms [110, 70].
Accelerating Attention Mechanisms.
The attention mechanism, with its quadratic computational complexity concerning context length, encounters increasing challenges as sequence lengths grow in modern large language models [101, 4, 5, 68, 103]. To address this limitation, polynomial kernel approximation methods [1] have been introduced, leveraging low-rank approximations to efficiently approximate the attention matrix. These methods significantly enhance computation speed, allowing a single attention layer to perform both training and inference with nearly linear time complexity [8, 10]. Moreover, these techniques can be extended to advanced attention mechanisms, such as tensor attention, while retaining almost linear time complexity for both training and inference [11][59] provides an almost linear time algorithm to accelerate the inference of VAR Transformer. Other innovations include RoPE-based attention mechanisms [9, 20] and differentially private cross-attention approaches [85]. Alternative strategies, such as the conv-basis method proposed in [70], present additional opportunities to accelerate attention computations, offering complementary solutions to this critical bottleneck. Additionally, various studies explore pruning-based methods to expedite attention mechanisms [73, 26, 74, 107, 106, 57, 119, 121].
Gradient Approximation.
The low-rank approximation is a widely utilized approach for optimizing transformer training by reducing computational complexity [82, 86, 10, 56, 26, 83]. Building on the low-rank framework introduced in [8], which initially focused on forward attention computation, [10] extends this method to approximate attention gradients, effectively lowering the computational cost of gradient calculations. The study in [82] further expands this low-rank gradient approximation to multi-layer transformers, showing that backward computations in such architectures can achieve nearly linear time complexity. Additionally, [86] generalizes the approach of [10] to tensor-based attention models, utilizing forward computation results from [11] to enable efficient training of tensorized attention mechanisms. Lastly, [56] applies low-rank approximation techniques during the training of Diffusion Transformers (DiTs), demonstrating the adaptability of these methods across various transformer-based architectures.